OpenStudy (anonymous):

At best buy, Larry buys three CDs and three DVDs for $96 from the sale rack. Kelly buys two CDs and one DVD from the same rack for $44. What is the price per CD and the price per DVD?

OpenStudy (anonymous):

First things first - go ahead and change the problem into a mathematic notation. Let's call CDs x and DVDs y. 3x+3y = 96 2x+1y = 44 Then we just solve our way for each variable through this linear system!

OpenStudy (anonymous):

thanks I do not know if I am correct but when I solved I got the answer x=32

OpenStudy (anonymous):

It's a bit too high for x. Because then we'd have y = 0 in the first equation and 66 = 44 in the second. Let's look at the second equation first 2x+1y = 44 First things first, let's solve for y here, just because it's easy. Move 2x to the right hand side by subtracting it from each side... We now have y = 44 - 2x. Solve for y in the first equation, and you'll see a pattern appear in the two, that'll help solve it.

OpenStudy (anonymous):

x=12 and y=20

OpenStudy (anonymous):

Fantastic! Good job!